# MEIT: Multi-Modal Electrocardiogram Instruction Tuning on Large Language Models for Report Generation

### v3

@@ Abstract
Electrocardiogram (ECG) is the primary non-invasive diagnostic tool for monitoring cardiac conditions and is crucial in assisting clinicians. Recent studies have concentrated on classifying cardiac conditions using ECG data but have overlooked ECG report generation, which is time-consuming and requires clinical expertise. To automate ECG report generation and ensure its versatility, we propose the Multimodal ECG Instruction Tuning (MEIT) framework, the first attempt to tackle ECG report generation with LLMs and multimodal instructions. To facilitate future research, we establish a benchmark to evaluate MEIT with various LLMs backbones across two large-scale ECG datasets. Our approach uniquely aligns the representations of the ECG signal and the report, and we conduct extensive experiments to benchmark MEIT with nine open-source LLMs using more than 800,000 ECG reports. MEIT’s results underscore the superior performance of instruction-tuned LLMs, showcasing their proficiency in quality report generation, zero-shot capabilities, and resilience to signal perturbation. These findings emphasize the efficacy of our MEIT framework and its potential for real-world clinical application.
@@ 1 Introduction
Diagnosing heart diseases is of paramount importance, requiring the analysis of multimodal data sources, which include electrocardiogram (ECG) recordings and clinical reports. Cardiologists read and interpret these ECG recordings to manually generate comprehensive ECG reports for heart disease diagnosis, which is a complex and time-consuming process. Recently, AI models have been developed to facilitate ECG data analysis for the task of classification [1, 2, 3, 4]. Despite these efforts, the automatic generation of reports from ECG recordings still needs to be explored. Unlike other AI-empowered medical report generation applications (e.g., radiology reports), the primary challenge for ECG report generation stems from the distinct nature of ECG content. ECG reports, often comprising brief phrases that summarize signal patterns, contrast with detailed anatomical descriptions in radiology reports. The difference in the content and semantic interpretation between imaging and ECG data complicates the direct application of radiology-focused methods to ECG reports. Furthermore, there is still a lack of comprehensive benchmarks for evaluating the performance of ECG report generation.
To tackle these challenges, we introduce MEIT, a Multimodal ECG Instruction Tuning framework that extends the capabilities of LLMs in the cardiology context to generate ECG reports using ECG recordings and human instructions. Inspired by the versatility of LLMs [5, 6, 7, 8, 9] in handling diverse language tasks simultaneously, we develop a specialized instruction tuning process for ECG report generation. MEIT aligns human instructions with ECG recordings, enabling LLMs to generate professional-grade reports and exhibit zero-shot report generation capabilities under domain transfer scenarios across various continents and data collection devices. Specifically, leveraging publicly available ECG datasets, we construct a multimodal instruction dataset including ECG records, human instructions, and paired reports. Then, we propose an effective and efficient attention-based fusion method to integrate ECG and text representations in the latent space. This enables LLMs to understand ECG signals for report generation without introducing additional training parameters in the attention layer. In addition to the ECG report generation approach, we introduce a comprehensive benchmark for ECG report generation evaluation, utilizing two datasets with 20K and 800K ECG-report pairs, respectively, across three evaluation tasks: report generation quality, zero-shot learning across datasets, and robustness analysis in the face of ECG signal perturbation. Utilizing the ECG report evaluation benchmark, we assess the proposed approach across nine open-source LLMs. The results demonstrate (1) the superior performance of MEIT in ECG report generation and the effective learning and alignment of ECG representations; (2) the effective transferability of LLMs under the MEIT framework in domain transfer scenarios.
To summarize, our primary contribution is the MELT framework, a novel approach to automating ECG report generation and evaluation based on LLMs. This framework incorporates a lightweight, attention-based fusion module across various LLM models. Furthermore, we design a new benchmark for ECG report generation, which contains three evaluation tasks. Our evaluations showcase the enhanced capabilities of instruction-tuned LLMs in generating ECG reports, highlighting the transferability in zero-shot tests and robustness against data perturbations. MEIT paves the way for future advances in automated ECG report generation and methodological innovations in integrating biomedical signals into LLMs.
2 Preliminaries
Electrocardiogram (ECG) measures the electrical activity of an individual’s heart over time. An ECG recording typically contains a 12-lead multivariate time series, which acts as a 12-dimensional sequence of embeddings. The ECG signal offers a comprehensive view, encompassing both spatial and temporal aspects of cardiac function. ECG leads can be categorized into six limb leads (i.e., I, II, III, aVR, aVL, and aVF) to monitor arms and legs, providing frontal plane views, and six precordial leads (i.e., V1, V2, V3, V4, V5, and V6) to monitor chest, showing horizontal plane views. We denote an ECG recording as $\mathbf{X}_{e}\in\mathcal{R}^{M\times T}$, where $M$ represents the number of leads, and $T$ is signal length. Each ECG recording is associated with an ECG report $\mathbf{X}_{t}$ for description. Thus, we denote each ECG pair as $\{\mathbf{X}_{e},\mathbf{X}_{t}\}$. More details on visualization can be found in the appendix A.5.
Figure 1: Illustration of the proposed MEIT framework, including ECG instruction data curation, instruction tuning on auto-regressive objective, and inference to get the generated report.
Figure 2: Illustration of model architecture for ECG Report Generation. $\mathbf{K}_{e}$ and $\mathbf{V}_{e}$ refer to linear projection of $\mathbf{H}_{e}$ by multiplying shared $\mathbf{W}_{k}$ and $\mathbf{W}_{v}$ in the attention layer.
@@ 3 MEIT
Figure 2 illustrates the proposed MEIT framework. We begin by describing the curation process for ECG instruction data. We then delve into the architecture of the multimodal LLM designed for ECG report generation. Lastly, we describe the approach to generating ECG reports with instruction tuning.
3.1 ECG Instruction Data Curation
@@
Given an ECG signal $\mathbf{X}_{e}$, our goal during inference is to generate an ECG report using an instruction prompt. For instance, the prompt can be “Given the ECG signal embeddings, please help me generate an accurate description for this ECG signal embeddings: ". To achieve this goal, during the training phase, we aim to create instruction tuning data to generate a response $\mathbf{\hat{X}}_{t}$ that aligns semantically with the ground truth $\mathbf{X}_{t}$. In addition, since we cannot predict the exact instruction prompt that users will use, we need to ensure that our report generation process is robust enough to handle different prompts. To address this challenge, we manually design some prompt samples, then utilize GPT-4 [5] to generate a set of prompts by rephrasing, as shown in Figure 2. Then, we randomly select one instruction prompt $\mathbf{X}_{p}$ from the prompt set and create a general instruction-following template: <|user|>: $\mathbf{X}_{p}$, $\mathbf{X}_{e}$}<|assistant|>: $\mathbf{X}_{t}$} </s>, where <|user|> and <|assistant|> are added special tokens for tokenizer, </s> is a stop sign for each response. This approach ensures that the generated response conveys the same meaning as the ground truth and remains adaptable to different instruction prompts. Following this strategy, we construct the ECG instruction data using a MIMIC-IV-ECG [10] dataset that contains 80K annotated data and a 20K dataset PTB-XL [11]. The ECG instruction data samples are shown in Appendix A.5.
3.2 ECG Report Generation Model
In MEIT, the multimodal ECG report generation model decodes ECG signals end-to-end to generate ECG reports. The architecture is illustrated in Figure 2. Specifically, the report generation model directly encodes an entire ECG-signal $\mathbf{X}_{e}\in\mathcal{R}^{M\times T}$ into latent embeddings and integrate it with the language embeddings with modality alignment, and then autoregressively generate the ECG report. Next, we detail each component of the report generation model.
@@ 3.2 ECG Report Generation Model
ECG Encoder. Since the ECG signal is of high resolution in the temporal domain, it is vital to efficiently extract temporal features per lead before interaction with semantic embeddings inside the LLM backbone. Our default ECG encoder $\mathcal{F}_{e}(\cdot)$ consists of temporal convolution blocks to encode each ECG signal into embeddings. Specifically, each temporal convolution block comprises several 1-D convolution layers, batch normalization layers, and ReLU activation layers, followed by average pooling. To align the output dimension with the head dimension of LLM backbone $\mathcal{F}_{l}(\cdot)$, we employ a non-linear projection layer $\mathcal{P}_{e}(\cdot)$ to obtain the ECG embeddings:
@@
where $\mathbf{H}_{e}\in\mathcal{R}^{D_{h}}$, $D_{h}$ has the same dimension as the multi-head attention layers of LLMs. Note that our default ECG encoder is lightweight and is able to learn temporal patterns of signals without a long training period. More details about ECG Encoder are illustrated in Appendix A.3.
ECG Modality Alignment. We introduce an ECG modality alignment strategy to guide the LLMs in aligning ECG signal data with corresponding textual outputs. This approach is detailed in Figure 2. Specifically, given the ECG embeddings $\mathbf{H}_{e}$, the alignment strategy incorporates $\mathbf{H}_{e}$ with the current hidden state $\mathbf{H}_{t}^{i}$ generated from previous $i-1$-th layer of the LLM backbone $\mathcal{F}_{l}\left(\cdot\right)$ for next-token prediction task. Here $\mathbf{H}_{t}^{i}$ is defined as:
@@
where $i$ is the current layer index. Different from gated-attention fusion in Flamingo [12], Memorizing Transformer [13] , and G-MAP [14], or Q-former in BLIP-2 [15] that require the additional trainable parameters and design complex modules for two-stage training, we propose a simple yet effective concatenated-fusion method that directly injects the ECG embeddings with language context in the causal-attention (C-Attn) of the models. In our approach, each attention layer combines $\mathbf{H}_{e}$, generated from the ECG encoder and projector as a prefix condition, with $\mathbf{H}_{t}^{i}$, derived from the preceding layer. The fusion process is as follows:
$\operatorname{C-Attn}\left(\mathbf{H}_{e},\mathbf{H}_{t}^{i}\right)=\left[\operatorname{head}_{1},\ldots,\operatorname{head}_{k}\right]\mathbf{W}_{o},$ (3)
@@
where $\mathbf{Q}_{t,j}=\mathbf{H}_{t,j}^{i}\mathbf{W}_{q,j}$, $\mathbf{K}_{e,j}=\mathbf{H}_{e}\mathbf{W}_{k,j}$, and $\mathbf{K}_{t,j}=\mathbf{H}_{t,j}^{i}\mathbf{W}_{k,j}$, with a similar notation for $\mathbf{V}_{e,j}=\mathbf{H}_{e}\mathbf{W}_{v,j}$ and $\mathbf{V}_{t,j}=\mathbf{H}_{t}^{i}\mathbf{W}_{v,j}$. Concatenation is denoted by []delimited-[][$\cdot$][ ⋅ ], and $\mathbf{K}_{m,j}$ and $\mathbf{V}_{m,j}$ symbolize the amalgamated features of query and key. $\mathbf{W}_{q,j}$, $\mathbf{W}_{k,j}$, and $\mathbf{W}_{v,j}$ in $\mathcal{R}^{D_{h}\times D_{h}}$ represent the projection matrices for query, key, and value for each head $j$, respectively. Our model’s design allows for the efficient fusion of two modalities through causal attention, facilitating conditional generation without the need for additional parameter updates to align the ECG modality. More comparisons about ECG modality alignment and other fusion approaches are illustrated in Appendix LABEL:appendix:_ECG_align_ablation.
3.3 ECG Instruction Tuning
As described in Section 3.1, we have converted ECG-text pairs into a chat-bot style instruction format: <|user|>: $\mathbf{X}_{p}$, $\mathbf{X}_{e}$}<|assistant|>: $\mathbf{X}_{t}$}</s>. During instruction tuning, we compute autoregressive loss only on tokens after response tokens <assistant>, and use label loss masking to finetune the model, where we mask all tokens belonging to $\mathbf{X}_{p}$ and $\mathbf{X}_{e}$. To save computational resources and accelerate the convergence of instruction tuning, we use LoRA [16] adapters for all linear layers of the LLM backbone $\mathcal{F}_{l}$ and freeze its backbone. Subsequently, given a sequence of ECG instruction data, we compute the probability of the target response $\mathbf{X_{t}}$ as an autoregressive function:

### v4

# MEIT: Multimodal Electrocardiogram Instruction Tuning on Large Language Models for Report Generation

@@ Abstract
Electrocardiogram (ECG) is the primary non-invasive diagnostic tool for monitoring cardiac conditions and is crucial in assisting clinicians. Recent studies have concentrated on classifying cardiac conditions using ECG data but have overlooked ECG report generation, which is time-consuming and requires clinical expertise. To automate ECG report generation and ensure its versatility, we propose the Multimodal ECG Instruction Tuning (MEIT) framework, the first attempt to tackle ECG report generation with LLMs and multimodal instructions. To facilitate future research, we establish a benchmark to evaluate MEIT with various LLMs backbones across two large-scale ECG datasets. Our approach uniquely aligns the representations of the ECG signal and the report, and we conduct extensive experiments to benchmark MEIT with nine open-source LLMs using more than 800,000 ECG reports. MEIT’s results underscore the superior performance of instruction-tuned LLMs, showcasing their proficiency in quality report generation, zero-shot capabilities, resilience to signal perturbation, and alignment with human expert evaluation. These findings emphasize the efficacy of MEIT and its potential for real-world clinical application. Our code is released at https://github.com/AIoT-MLSys-Lab/MEIT.
MEIT: Multimodal Electrocardiogram Instruction Tuning on
Large Language Models for Report Generation
Zhongwei Wan1, Che Liu2, Xin Wang1, Chaofan Tao3, Hui Shen1, Jing Xiong3, Rossella Arcucci2, Huaxiu Yao4, Mi Zhang1 1The Ohio State University  2 Imperial College London 3The University of Hong Kong  4University of North Carolina at Chapel Hill https://github.com/AIoT-MLSys-Lab/MEIT
@@ 1 Introduction
Electrocardiogram (ECG) is the primary tool for heart disease diagnosis. In standard practice, cardiologists examine these ECG recordings and manually generate detailed diagnostic reports, which is a complex and time-consuming process. Recently, AI models have been developed to streamline ECG data analysis for the classification task (Hu et al., 2023; Liu et al., 2023a, 2024a), yet the automatic generation of reports from ECG recordings remains relatively underexplored.
Unlike image-based medical report generation tasks (e.g., radiology reports), ECG report generation faces unique challenges due to the concise, keyword-centric content of ECG reports, which contrasts with the more extensive anatomical descriptions in radiology. Directly transferring methods from radiology is hindered by the distinct nature of ECG signals and the limited semantic overlap with imaging data. Furthermore, there is still a lack of comprehensive benchmarks for evaluating the performance of ECG report generation.
Drawing on the versatility of LLMs and multimodal LLMs (MLLMs) (Achiam et al., 2023; Touvron et al., 2023b; Wan et al., 2023b; Wang et al., 2024a; Zhu et al., 2024b; Wang et al., 2024b; Zhang et al., 2025a; Zhu et al., 2025) in various language tasks, in this work, we introduce MEIT, a Multimodal ECG Instruction Tuning framework that extends the capability of large language models (LLMs) for the task of ECG report generation. By aligning ECG recordings with human instructions, MEIT produces clinically relevant reports and demonstrates zero-shot performance when transferring across different continents and data collection devices. Concretely, we construct a multimodal instruction dataset from publicly available ECG recordings and propose an efficient attention-based fusion method that incorporates ECG signals into the latent space of LLMs without adding new parameters.
We also introduce a comprehensive benchmark to evaluate ECG report generation across two large-scale datasets (with 20K and 800K ECG-report pairs), covering four tasks: (1) report generation quality, (2) zero-shot transfer across datasets, (3) robustness under signal perturbations, and (4) alignment with expert evaluation. We assess MEIT using ten open-source LLMs, demonstrating: (1) the superior performance of MEIT in ECG report generation and effective ECG representation learning; and (2) the strong transferability of instruction-tuned LLMs across diverse clinical domains.
In summary, our primary contribution is the MEIT framework, a novel method for automated ECG report generation and evaluation based on LLMs. It features a lightweight, attention-based fusion module for integrating ECG signals and text, along with a newly designed four-task benchmark for ECG report evaluation. Empirical results highlight MEIT’s ability to generate high-quality reports, perform robustly under data perturbations, and align closely with expert assessments, thereby paving the way for improved ECG interpretation and broader innovations in embedding biomedical signals into LLMs.
2 Related Work
Medical Report Generation. Our work is closely related to medical report generation, which has been extensively studied in the context of medical images. Existing methods can be grouped into three main categories: (1) template-based methods, such as HRGR (Li et al., 2018) and CMAS (Jing et al., 2017); (2) data integration and coherence methods, such as PPKED (Liu et al., 2021) and CA (Ma et al., 2021); and (3) cross-modal alignment methods, exemplified by Chen et al. (2022) and Qin and Song (2022). However, these methods are designed for medical images and struggle with ECG signals due to their distinct temporal and waveform characteristics.
Instruction Tuning. Our work is also related to instruction tuning, which improves zero-shot learning in LLMs through task-specific instructions (Zhang et al., 2023; Wang et al., 2023). Models like InstructGPT (Ouyang et al., 2022), FLAN-PaLM (Chung et al., 2022), and Alpaca (Taori et al., 2023) leverage instruction data, including human feedback. Multimodal variants, such as LLaVA (Liu et al., 2023c), MiniGPT-4 (Zhu et al., 2023), and AnyMAL (Moon et al., 2023), extend this to visual tasks. However, these methods focus on natural images and are not directly applicable to ECG signals. To bridge this gap, we introduce a specialized instruction-tuning framework for ECG report generation.
LLMs for ECG. Few studies have explored the use of LLM for the analysis of ECG signals (Liu et al., 2023b; Qiu et al., 2023; Yu et al., 2023). Some of them (Liu et al., 2023b; Yu et al., 2023) convert ECG signals into text features before importing them into LLMs, bypassing crucial modality-specific characteristics. In addition, these works focus on the task of disease classification rather than report generation. In contrast, our method directly processes ECG signals and focuses on the task of report generation.
Figure 1: (a) Overview of MEIT; (b) Illustration of the architecture of Report Generator.
@@ 3 MEIT
3.1 Preliminaries
Electrocardiogram (ECG) measures the electrical activity of an individual’s heart over time. An 12-lead ECG recording is a multivariate time series. It offers a multi-dimensional view, encompassing both spatial and temporal aspects of cardiac function. The 12 ECG leads include six limb leads (i.e., I, II, III, aVR, aVL, and aVF) that monitor arms and legs, providing frontal plane views, and six precordial leads (i.e., V1, V2, V3, V4, V5, and V6) that monitor chest, showing horizontal plane views.
Formally, let $\mathbf{X}_{e}\in\mathcal{R}^{M\times T}$ denote an ECG recording, where $M$ represents the number of leads, and $T$ represents the signal length. Each ECG recording is associated with an ECG report $\mathbf{X}_{t}$ that describes and interprets the ECG recording. Thus, we denote each ECG recording-report pair as $\{\mathbf{X}_{e},\mathbf{X}_{t}\}$.
3.2 Overview of MEIT
Figure 1 (a) illustrates the proposed MEIT framework. In the data curation stage, we construct the ECG instruction tuning data, which includes instruction prompts, ECG recordings and the corresponding ground truth ECG reports. In the instruction tuning stage, the ECG instruction tuning data are fed into the Report Generator for training using an autoregressive approach. During inference, the instruction prompts and the ECG recordings are inputted into the Report Generator to generate the ECG reports. In the following sections, we describe each component in detail.
3.3 Data Curation
@@
Given an ECG recording $\mathbf{X}_{e}$, our goal during inference is to generate an ECG report using an instruction prompt such as “Given the ECG recording, please help me generate a report for this ECG recording:". To achieve this goal, we aim to create a set of instruction tuning data to generate a response $\mathbf{\hat{X}}_{t}$ that semantically aligns with the ground truth $\mathbf{X}_{t}$. Since we cannot predict the exact instruction prompt that users will use, we need to ensure that our report generation process is robust enough to handle different prompts. To do so, we manually design a small set of prompt samples, and then utilize GPT-4 (Achiam et al., 2023) to generate a larger prompt set by rephrasing, as shown in Figure 1 (a). For each ECG recording-report pair as $\{\mathbf{X}_{e},\mathbf{X}_{t}\}$, we randomly select one instruction prompt $\mathbf{X}_{p}$ from the larger prompt set and create an instruction tuning template: <|user|>: $\mathbf{X}_{p}$, $\mathbf{X}_{e}$} <|assistant|>: $\mathbf{X}_{t}$} </s>, where <|user|> and <|assistant|> are added special tokens for tokenizer, and </s> is a stop sign for each response. This approach ensures that the generated response conveys the same meaning as the ground truth and remains adaptable to different instruction prompts. Following this strategy, we construct the ECG instruction tuning data using the PTB-XL (Wagner et al., 2020) dataset and the MIMIC-IV-ECG (Gow et al., ) dataset. Some examples of ECG instruction tuning data are included in Appendix A.8.
3.4 Report Generation
Figure 1 (b) illustrates the architecture of the Report Generator. As shown, the Report Generator utilizes an ECG encoder to encode $\mathbf{X}_{e}$ into ECG embeddings and integrates them with the language embeddings extracted from the corresponding instruction tuning template through an ECG-text alignment module to autoregressively generate the ECG report. In this section, we describe the key components of the Report Generator in detail.
@@ 3.4 Report Generation
ECG Encoder. Since the ECG recording is high resolution in the time domain, it is vital to efficiently extract temporal features per lead before interaction with semantic embeddings inside the LLM backbone. Our default ECG encoder $\mathcal{F}_{e}(\cdot)$ consists of temporal convolution blocks to encode the ECG recordings into embeddings. Specifically, each temporal convolution block comprises several 1-D convolution layers, batch normalization layers, and ReLU activation layers, followed by average pooling. This design allows us to effectively capture temporal dependencies and reduce the complexity of the signal representations, ensuring that the model can quickly learn important temporal features efficiently. To further align the output dimension with the head dimension of the LLM backbone $\mathcal{F}_{l}(\cdot)$, we employ a non-linear projection layer $\mathcal{P}_{e}(\cdot)$ to generate the ECG embeddings:
@@
where $\mathbf{H}_{e}\in\mathcal{R}^{D_{h}}$, $D_{h}$ has the same dimension as the multi-head attention layers of LLMs. Note that our default ECG encoder is lightweight and is able to learn temporal patterns of ECG recordings without a long training period. More details about the ECG encoder are included in Appendix A.3.
ECG-Text Alignment. Given the ECG and the text embeddings, we introduce an ECG-text alignment strategy to guide the LLM in aligning the ECG embeddings with the text embeddings. As shown in Figure 1 (b), given the ECG embeddings $\mathbf{H}_{e}$, the ECG-text alignment strategy incorporates $\mathbf{H}_{e}$ with the current hidden state $\mathbf{H}_{t}^{i}$ generated from the previous ($i-1$)th layer of the LLM backbone $\mathcal{F}_{l}\left(\cdot\right)$ for next-token prediction task. Here $\mathbf{H}_{t}^{i}$ is defined as:
@@
where $i$ is the current layer index.
Traditional gated-attention fusion methods like Flamingo (Alayrac et al., 2022), Memorizing Transformer (Wu et al., 2022), G-MAP (Wan et al., 2022), and Q-former in BLIP-2 (Li et al., 2023) require additional trainable parameters and are designed for complex multi-stage alignment of rich semantic information (e.g., images). Different from them, our method provides a lightweight concatenation-based alignment strategy tailored to the ECG embeddings, enabling efficient learning of ECG semantic features via directly injecting the ECG embeddings with language context in the self-attention, while preventing potential catastrophic forgetting of general knowledge in LLMs.
In our approach, each attention layer combines $\mathbf{H}_{e}$, generated from the ECG encoder and projector as a prefix condition, with $\mathbf{H}_{t}^{i}$, derived from the preceding layer. The fusion process is as follows:
$\operatorname{Self-Attn}\left(\mathbf{H}_{e},\mathbf{H}_{t}^{i}\right)=\left[\operatorname{head}_{1},\ldots,\operatorname{head}_{k}\right]\mathbf{W}_{o},$ (3)
@@
where $\mathbf{Q}_{t,j}=\mathbf{H}_{t,j}^{i}\mathbf{W}_{q,j}$, $\mathbf{K}_{e,j}=\mathbf{H}_{e}\mathbf{W}_{k,j}$, and $\mathbf{K}_{t,j}=\mathbf{H}_{t,j}^{i}\mathbf{W}_{k,j}$, with a similar notation for $\mathbf{V}_{e,j}=\mathbf{H}_{e}\mathbf{W}_{v,j}$ and $\mathbf{V}_{t,j}=\mathbf{H}_{t}^{i}\mathbf{W}_{v,j}$. Concatenation is denoted by []delimited-[][$\cdot$][ ⋅ ], and $\mathbf{K}_{m,j}$ and $\mathbf{V}_{m,j}$ symbolize the amalgamated features of query and key. $\mathbf{W}_{q,j}$, $\mathbf{W}_{k,j}$, and $\mathbf{W}_{v,j}$ in $\mathcal{R}^{D_{h}\times D_{h}}$ represent the projection matrices for query, key, and value for each head $j$, respectively. Such design allows for the efficient fusion of two modalities through causal attention, facilitating conditional generation without the need for additional parameter updates to align the ECG modality with the text modality. Ablation studies comparing with other fusion methods demonstrate the effectiveness and efficiency of our proposed lightweight alignment strategy. More comparisons about ECG-text alignment and other fusion approaches are included in Table 5.
3.5 Instruction Tuning
During instruction tuning, we follow the instruction tuning template: <|user|>: $\mathbf{X}_{p}$, $\mathbf{X}_{e}$} <|assistant|>: $\mathbf{X}_{t}$} </s> and compute the autoregressive loss only on tokens after response tokens <assistant>, and use label loss masking to finetune the model, where we mask all tokens belonging to $\mathbf{X}_{p}$ and $\mathbf{X}_{e}$. To save computational resources and accelerate the convergence of instruction tuning, we use LoRA (Hu et al., 2021) adapters for all linear layers of the LLM backbone $\mathcal{F}_{l}$ and freeze its backbone. Subsequently, given a sequence of ECG instruction data, we compute the probability of the target response $\mathbf{X_{t}}$ as an autoregressive function: